40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Feb 20 23:55:51 2019, gautam, Update, ALS, ALS delay line electronics delayLineDemod.pdf.zip
    Reply  Thu Mar 7 01:06:38 2019, gautam, Update, ALS, ALS delay line electronics 
       Reply  Tue Mar 12 22:51:25 2019, gautam, Update, ALS, ALS delay line electronics 
       Reply  Wed Mar 13 01:27:30 2019, gautam, Update, ALS, ALS delay line electronics DFDcal.pdfDFDnoise.pdf
          Reply  Sun Mar 17 00:42:20 2019, gautam, Update, ALS, NF1611 cannot be shot-noise limited? PDnoise.pdf
             Reply  Thu Mar 28 19:40:02 2019, gautam, Update, ALS, BeatMouth with NF1611s assembled IMG_7381.JPGIMG_7382.JPGrelTF_schem.pdfdarkNoise.pdf
                Reply  Fri Mar 29 21:00:06 2019, gautam, Update, ALS, BeatMouth with NF1611s installed IMG_7384.JPGIMG_7385.JPG
                   Reply  Sun Mar 31 15:05:53 2019, gautam, Update, ALS, Fiber beam-splitters not PM IMG_7384.JPG
                      Reply  Wed Apr 3 09:04:01 2019, gautam, Update, ALS, Note about new fiber couplers 
                         Reply  Wed Apr 3 12:32:33 2019, Koji, Update, ALS, Note about new fiber couplers 
                      Reply  Fri Apr 5 00:33:58 2019, gautam, Update, ALS, Promising IR ALS noise IMG_7388.JPGfreeSpace_IR_beat.pdf
                         Reply  Fri Apr 5 11:49:30 2019, gautam, Update, ALS, PSL + X green beat recovery 
                         Reply  Mon Apr 8 00:04:08 2019, gautam, Update, ALS, IR ALS noise budget DFDnoise.pdf
                            Reply  Mon Apr 8 18:28:25 2019, gautam, Update, ALS, EX Green PDH checkout IMG_7393.JPGgreenModeMatch.pdf
                               Reply  Mon Apr 8 23:52:09 2019, gautam, Update, ALS, EX Green PDH error monitor calibrated errMonCalib.pdferrMon.pngOLTF.pdfEX_frequencyNoises.pdf
                                  Reply  Tue Apr 9 00:16:22 2019, rana, Update, ALS, EX Green PDH error monitor calibrated 
                                     Reply  Tue Apr 9 00:18:19 2019, gautam, Update, ALS, EX Green PDH error monitor calibrated 
                                  Reply  Tue Apr 9 18:44:00 2019, gautam, Update, ALS, EX Green PDH discriminant measurement cavityFlashes.pdfmodelPDH.pdf
                            Reply  Tue Apr 9 19:07:12 2019, gautam, Update, ALS, IR ALS noise budget ALS_noiseBudget.pdfALS_noiseBudget.pdf
                               Reply  Thu Apr 11 01:10:05 2019, gautam, Update, ALS, Large 2kHz peak (and harmonics) in ALS X EX_PDH_2kNoise.pdf
                                  Reply  Wed Apr 17 00:50:17 2019, gautam, Update, ALS, Large 2kHz peak (and harmonics) in ALS X no more EX_PDHnoise.pdf
                                     Reply  Wed Apr 17 11:01:49 2019, gautam, Update, ALS, Large 2kHz peak (and harmonics) in ALS X no more EX_PDH_specGram.pdf
                      Reply  Wed May 29 18:13:25 2019, gautam, Update, ALS, Fiber beam-splitters are now PM 
                         Reply  Fri May 31 15:55:16 2019, gautam, Update, ALS, PSL + X beat restored 
Message ID: 14498     Entry time: Thu Mar 28 19:40:02 2019     In reply to: 14480     Reply to this: 14502
Author: gautam 
Type: Update 
Category: ALS 
Subject: BeatMouth with NF1611s assembled 

Summary:

The parts I was waiting for arrived. I finished the beat mouth assembly, and did some characterization. Everything looks to be working as expected.

Details:

Attachment #1: Photo of the front panel. I am short of two fiber mating sleeves that are compatible with PM fibers, but those are just for monitoring, so not critical to the assembly at this stage. I'll ask Chub to procure these.

Attachment #2: Photo of the inside of the BeatMouth. I opted to use the flexible RG-316 cables for all the RF interconnects. Rana said these aren't the best option, remains to be seen if interference between cables is an issue. If so, we can replace them with RG-58. I took the opportunity to give each fiber beam splitter its own spool, and cleaned all the fiber tips.

Attachment #3: Transfer function measurement. The PDFR setup behind 1X5/1X6 was used. I set the DC current to the laser to 30.0 mA (as read off the display of the current source), which produced ~400uW of light at the fiber coupled output of the diode laser. This was injected into the "PSL" input coupler of the BeatMouth, and so gets divided down to ~100 uW by the time it reaches the PDs. From the DC monitor values (~430mV), the light hitting the PDs is actually more consistent with 60uW, which is in agreement with the insertion loss of the fiber beamsplitters, and the mating sleeves.

The two responses seem reasonably well balanced (to within 20% - do we expect this to be better?). Even though judging by the DC monitor, there was more light incident on the Y PD than on the X PD, the X response was actually stronger than the Y. 

I also took the chance to do some other tests:

  • Inject light into the "X(Y)-ARM" input coupler of the Beat Mouth - confirmed that only the X(Y) NF1611's DC monitor output showed any change. The DC light level was ~1V in this condition, which again is consistent with expected insertion losses as compared to the "PSL" input case, there is 1 less fiber beamsplitter and mating sleeve.
  • Injected light into each of the input couplers, looked at the interior of the BeatMouth with an IR viewer for evidence of fiber damage, and saw none. Note that we are not doing anything special to dump the light at the unused leg of the fiber beamsplitter (which will eventually be a monitor port). Perhaps, nominally, this port should be dumped in some appropriate way.

Attachment #4: Dark Noise analysis. I used a ZHL-500-HLN+ to boost the PD's dark noise above the AG4395's measurement noise floor. The measured noise level seems to suggest either (i) the input-referred current noise of the PD circuitry is a little lower than the spec of 16 pA/rtHz (more like 13 pA/rtHz) or (ii) the transimpedance is lower than the spec of 700 V/A (more like 600 V/A). Probably some combination of the two. Seems reasonable to me.

Next steps:

The optical part of the ALS detection setup is now complete. The next step is to measure the ALS noise with this sysytem. I will use the X arm for this purpose (I'd like to make the minor change of switching the existing resistive power splitter at the delay line to the newly acquired splitters which have 3dB lower insertion loss). 

Attachment 1: IMG_7381.JPG  1.628 MB  Uploaded Thu Mar 28 23:53:25 2019  | Hide | Hide all
IMG_7381.JPG
Attachment 2: IMG_7382.JPG  1.979 MB  Uploaded Fri Mar 29 00:04:21 2019  | Hide | Hide all
IMG_7382.JPG
Attachment 3: relTF_schem.pdf  609 kB  Uploaded Fri Mar 29 00:24:58 2019  | Hide | Hide all
relTF_schem.pdf
Attachment 4: darkNoise.pdf  55 kB  Uploaded Fri Mar 29 13:33:39 2019  | Hide | Hide all
darkNoise.pdf
ELOG V3.1.3-